# pQT_AMTs.bib

### v1

@@ Abstract
Are there consistent and physically reasonable alternatives to the projection postulate? Does it have unique properties compared to acceptable alternatives? We answer these questions by systematically investigating hypothetical state-update rules for quantum systems that nature could have chosen over the Lüders rule. Among other basic properties, any prospective rule must define unique post-measurement states and not allow for superluminal signalling. Particular attention will be paid to consistently defining post-measurement states when performing local measurements in composite systems. Explicit examples of valid unconventional update rules are presented, each of them resulting in a distinct, well-defined foil of quantum theory. This framework of state-update rules allows us to identify operational properties that distinguish the projective update rule from all others and to put earlier derivations of the projection postulate into perspective.
@@ 1 Introduction
The main goal of this contribution is to identify operational properties that distinguish the quantum-mechanical projection postulate from other physically reasonable, but hypothetical, state-update rules. Justifications of Lüders rule have, not surprisingly, a long history, and a recent debate revolves around the claim that it can be derived from the other postulates of quantum theory (cf. Sec. 2.1). Many of these approaches have in common that their focus is on non-composite systems, either ignoring measurements on composite systems or making implicit assumptions about their consequences.
We study the nature of the projection postulate systematically by creating foil theories closely resembling quantum theory but equipped with different generalised state-update rules (GURs). In other words, we introduce generalised state-update theories (GUTs) for which the post-measurement states associated with a given outcome are not necessarily those prescribed by the Lüders rule. From the outset, we will take into account both single and composite systems. Alternative post-measurement rules have been considered occasionally (cf. Sec. 2.1), but we are not aware of any systematic approach.
@@ 2.1 The quantum-mechanical state update rule
Using quantum systems as information carriers has led to applications without classical equivalents. Quantum measurements, in particular, provide a resource that contributes, for example, to secure communication protocols [bennett_quantum_2014], quantum teleportation [bennett_teleporting_1993] device-independent random number generators [colbeck_quantum_2009] and to speedy algorithms when using measurement-based quantum computation [briegel_measurement-based_2009].
The advantages of quantum measurements arise from fundamental differences when compared to measurements in classical theories. Classical measurements are deterministic in that they reveal pre-existing values of observables, and the effect on the state of the observed system is negligible, at least in principle. In quantum theory, measurement outcomes are probabilistic, and the state of a system after a measurement coincides with the pre-measurement state only in exceptional cases. Phenomenologically, the quantum mechanical state update is governed by the projection postulate, which describes a “collapse” of the state to an eigenstate of the operator representing the observed outcome. The collapse, also known as state reduction or Lüders rule [luders_uber_1950], ensures deterministic repeatability: two measurements of an observable always yield identical outcomes when implemented in quick succession [compton_directed_1925].
The standard postulates of quantum theory neither explain whether the state-update is a physical process [bohm_suggested_1952, dewitt_many-worlds_2015], nor do they define what constitutes a measurement [bell_against_1990, bacciagaluppi_heisenberg_2009]. The foundational questions resulting from this dilemma are often grouped under the term “measurement problem[busch_quantum_1996, norsen_foundations_2017]. The present work does not attempt to resolve the interpretational challenges posed by this long-standing problem.
Justifications of the Lüders rule have been given [bell_moral_1966, herbut_derivation_1969, hellmann_quantum_2016, kostecki_luders_2014, martinez_search_1990] (see Appendix A for a survey), and the difference between von Neumann’s and Lüders’ formulations of the projection postulate has been discussed repeatedly [vonneumann_mathematische_1932, luders_uber_1950, busch_luders_1998, sudbery_whose_2024]. Some authors have argued that the projection postulate follows from the other standard postulates of quantum theory [ozawa_quantum_1984, masanes_measurement_2019]; however, these derivations—one of which was the subject of a recent debate [stacey_masanes-galley-muller_2022, galley_reply_2022, kent_measurements_2023, masanes_response_2023]—appear to invariably rely on additional assumptions.
Occasionally, alternative post-measurement rules have been considered, either in the context of standard quantum theory [kent_nonlinearity_2005, kent_quantum_2021, masanes_measurement_2019] or of generalised probabilistic theories [kleinmann_sequences_2014]. A systematic investigation, as proposed here, has apparently not been carried out before.
@@ 2.2 Foil theories
Our approach is inspired by the constructive role of foil theories for the understanding of quantum theory. The strategy in defining foil theories is to embed quantum theory in a larger set of theories by relaxing specific assumptions about the structure of the state space, say, or the form of the time evolution. We briefly describe typical examples of foil theories and point out that a systematic investigation of generalised state-update rules has, apparently, not yet been carried out.
@@
Generalised probabilistic theories (GPTs) [barrett_information_2007, hardy_quantum_2001, janotta_generalized_2014], for example, represent a large and versatile collection of foil theories. They are defined by their own sets of states, observables, and rules for system composition, all of which may differ from those of standard quantum theory. GPTs have become a standard tool to explore and identify distinctive characteristics of quantum theory which they contain as a special case. Many GPTs share properties such as the no-cloning theorem, uncertainty relations, or teleportation with quantum theory [barnum_cloning_2006, oppenheim_uncertainty_2010]. Some of them reproduce the standard quantum-mechanical correlations [wright_gleason-type_2019], while others exhibit superquantum correlations, e.g. Boxworld [popescu_quantum_1994].
@@
Ontological models describe a conceptually different class of foil theories not necessarily based on the familiar Hilbert space structure of quantum theory. They aim to explain its predictions in terms of an observer’s incomplete knowledge about the ontic state that provides an objective and complete description of the state of a system. In these models, non-classical features of quantum theory such as complementarity of observables, superdense coding, interference phenomena, or uncertainty relations may result from epistemic restrictions [spekkens_evidence_2007, bartlett_reconstruction_2012].
@@
Other foil theories differ from quantum theory in a single key feature only, such as attempts to modify Born’s rule [galvan_generalization_2008, galley_classification_2017]. The consequences of associating observables with non-Hermitian operators have been investigated, using PT-symmetric [bender_pt_2024] or normal [roberts_observables_2018] operators, for example. Quantum-like theories defined over the real numbers [stueckelberg_quantum_1960, renou_quantum_2021] or quaternions [finkelstein_foundations_1962] rather than the field of complex numbers have also been studied.
@@
Nonlinear generalisations of Schrödinger’s equation have been suggested [beretta_quantum_1984, abrams_nonlinear_1998, ferrero_nonlinear_2004, kent_nonlinearity_2005, rembielinski_nonlinear_2020], most notably in Weinberg’s work [weinberg_testing_1989]. Gisin et al. [gisin_weinbergs_1990, simon_no-signaling_2001, bassi_no-faster-than-light-signaling_2015] pointed out that they imply violations of the no-signalling principle, although some specific nonlinear transformations were later found not to be ruled out by Gisin’s argument [czachor_nonlocal_1998, kent_nonlinearity_2005, ferrero_nonlinear_2004, helou_extensions_2017, rembielinski_nonlinear_2020].
@@
In the specific context of modifications to the projection postulate, “causal quantum theory[kent_causal_2005] has been proposed, positing that the Lüders collapse is a well-defined physical process that satisfies strict local causality. Additionally, Kent introduced a new type of hypothetical measurement that is capable of revealing complete or partial information about the “local” state of a system, without actually inducing a state update [kent_nonlinearity_2005, kent_quantum_2021, kent_measurements_2023], an idea first presented in [busch_is_1997]. Such a measurement would, effectively, allow for yet other nonlinear state transformations without enabling superluminal signalling.
@@ 3.1 Axioms of quantum theory
    The probability that a measurement of $M$ performed on a system in the state $\rho$ yields the outcome represented by $P_{x}$ is given by the Born rule: $\text{prob}(x)=\text{Tr}\left[P_{x}^{A}\,\rho\right]$.
@@
which sends $\rho$ to a sub-normalized state [^1] A sub-normalised state is a non-negative operator on a Hilbert space with trace less than or equal to $1$, i.e. $\rho\in\bar{\mathcal{S}}(\mathcal{H})=\{\lambda\rho:\lambda\in[0,1],\rho\in\mathcal{S}(\mathcal{H})\}$. in the space $\bar{\mathcal{S}}(\mathcal{H})$, and call it the Lüders rule, giving postulate ($\mathsf{L}$) its name. The trace of the post-measurement state, $\text{Tr}(\omega^{\mathsf{L}}(P_{x},\rho))=\text{Tr}(P_{x}\rho)$, yields the outcome probability as prescribed by the Born rule of postulate ($\mathsf{B}$). In addition, the maps (1) and Eq. (2) do not require the input state to be normalised. Replacing $\rho$ by $\lambda\rho$, $\lambda\in[0,1]$, shows that the Lüders rule is 1-homogeneous: $\omega^{\mathsf{L}}(P_{x},\lambda\rho)=\lambda\omega^{\mathsf{L}}(P_{x},\rho)$.
The alternative update rules $\omega(P_{x},\rho)$ considered in Sec. 5 will be modeled on the Lüders rule in the sense that they are 1-homogeneous, only depend on the measurement outcome $P_{x}$ and the initial state $\rho$, and agree with the Born rule for outcome probabilities. Further physically motivated properties that any acceptable update rule should satisfy will be discussed in Sec. 4.
@@ 3.2 Measurements on composite quantum systems
There are, however, consistency requirements when performing measurements on a composite system. If the outcome $P_{x}^{A}$ is non-degenerate, then the Lüders rule assigns a pure post-measurement state $\omega^{\mathsf{L}}(P_{x}^{A},\rho_{A})$ to subsystem $A$. Consequently, the post-measurement state of the composite system must be a product state, whatever the initial state $\rho_{AB}$. On its own, the projection postulate ($\mathsf{L}$) provides, however, no information about the local state at $B$ upon performing a local measurement of $M_{A}$ on the constituent $A$. In principle, any density matrix $\rho_{B}=\text{Tr}_{B}(\rho_{AB})\in\mathcal{S}(\mathcal{H}_{B})$ is compatible. For example, one could imagine that any local measurement at $A$ throws the state of the composite system into a product state, with the post-measurement state at $B$ being equal to the maximally mixed state $\rho_{B}\propto\mathbb{I}_{B}$.
@@ 4.2 Necessary properties of generalised update rules
The first three assumptions do not explicitly rely on the composite structure. Completeness requires that every conceivable measurement outcome specifies a unique post-measurement state. Context-independence asserts that the observed outcome and the pre-measurement state alone determine the post-measurement state. It introduces a form of non-contextuality [kochen_problem_1967] that is present in quantum theory—namely, that all operationally equivalent experimental runs (i.e. involving the same initial state and the same measurement outcome) yield identical post-measurement states. Mathematically, completeness and context-independence determine the co-domain and the domain of the update rule $\omega_{AB}$ for a composite system, respectively.
@@
Completeness &\&& context-independence A complete and context-independent update rule maps any combination of a (local) outcome, $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, and a sub-normalised state, $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, to another sub-normalised state,
@@
A generalised state-update rule $\omega$ is a set of functions $\{\omega_{AB}\}$—where each function $\omega_{AB}$ is associated with a pair of Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$—satisfying the following properties: (i) consistency with the Born rule (Eq. (6)); (ii) 1-homogeneity (Eq. (7)); and (iii) assumptions A1-A6 (Eqs. (8)-(13)).
@@
How do the update rules defined here compare with quantum instruments [heinosaari_mathematical_2011], i.e. collections $\{\mathcal{I}_{x}\}_{x}$ of quantum operations [^7] Quantum operations are linear, completely positive and trace non-increasing maps defined on the space $\mathcal{L}(\mathcal{H})$ of bounded operators acting on $\mathcal{H}$. Quantum instruments are usually defined as mappings from an outcome space $\left(X,\Sigma\right)$ to the set of quantum operations. However, since we will only deal with discrete observables, an instrument is completely determined by the finite set of operations $\{\mathcal{I}_{x}\}_{x}$. that describe the effect on states of quantum measurements? The main difference is that the state-update rules satisfying our definition are not required to be linear or completely positive. Consequently, the set of alternative state-update rules is larger than the set of quantum instruments.
@@ 4.3 Updating (im-)proper mixtures
For a pure joint state $\rho_{AB}$, Eq. (11) applies irrespective of $\rho_{AB}$ being a product state or being entangled. Accordingly, $\omega_{A}\left(P_{x}^{A},\rho_{A}\right)$ represents the post-measurement state of system $\mathcal{H}_{A}$ when $\rho_{A}$ is either a pure state or an improper mixture, i.e. the reduced density operator of an entangled state. In contrast, proper mixtures result from classical uncertainty about the “true” preparation of a system. A proper mixture can be thought of as a Gemenge (German for ‘mixture’), i.e. as a collection $\mathcal{G}=\{(p_{1},\rho_{1}),(p_{2},\rho_{2}),\dots,(p_{n},\rho_{n})\}$ where $p_{i}\geq 0$, $\sum_{i}p_{i}=1$ and $\rho_{i}\in\bar{\mathcal{S}}(\mathcal{H})$ represent pure states or improper mixed states. The Gemenge $\mathcal{G}$ describes a system that is prepared in state $\rho_{i}$ with probability $p_{i}$ [busch_quantum_1996]. To any Gemenge there corresponds a unique density operator, $\rho_{\mathcal{G}}=\sum_{i=1}^{n}p_{i}\rho_{i}$.
@@
The equations (8)-(13) formalising assumptions A1-6 were formulated with pure and improper mixed states in mind. Nevertheless, they also ensure a consistent treatment of measurements on proper mixtures. For instance, the completeness A1 of $\omega$ guarantees that $\rho_{\mathcal{G}_{x}}$ in (14) is a valid state of the system. Similar arguments apply to assumptions A3-6, with the exception of context-independence A2, which requires minor adaptation. In fact, the post-measurement state $\rho_{\mathcal{G}_{x}}$ generally does not depend on the initial mixed state $\rho_{\mathcal{G}}$, but rather on the initial Gemenge $\mathcal{G}$. When proper mixtures are included, context-independence is generalised as follows: the observed outcome and the pre-measurement Gemenge alone determine the post-measurement state. [^9] If $\rho_{A}$ denotes either a pure or an improper mixed state, we can associate the trivial Gemenge $\mathcal{G}=\{(1,\rho_{A})\}$, describing a system prepared in the state $\rho_{A}$ with probability $1$.

### v2

# Beyond the Projection Postulate and Back: Quantum Theories with Generalised State-Update Rules

@@ Abstract
Are there consistent and physically reasonable alternatives to the projection postulate? Does it have unique properties compared with acceptable alternatives? We answer these questions by systematically investigating hypothetical state-update rules for quantum systems that Nature could have chosen over the Lüders rule. Among other basic properties, any prospective rule must define unique post-measurement states and not allow for superluminal signalling. Particular attention will be paid to consistently defining post-measurement states when performing local measurements in composite systems. Explicit examples of valid unconventional update rules are presented, each resulting in a distinct, well-defined foil of quantum theory. This framework of state-update rules allows us to identify operational properties that distinguish the projective update rule from all others and to put earlier derivations of the projection postulate into perspective.
@@ 1 Introduction
The main goal of our contribution is to identify operational properties that distinguish the quantum-mechanical projection postulate from hypothetical but physically reasonable state-update rules. Not surprisingly, attempts to justify Lüders rule (for which both Dirac and von Neumann deserve credit) [^2] In view of the historic developments, it would be appropriate to speak of the “Dirac–von Neumann-Lüders rule” [undefaay] but we will continue using “Lüders rule” for simplicity. However, the projection postulate as formulated by Dirac [undefac] and Lüders [undefaaf] differs from von Neumann’s version [undefaaz] in key aspects; see Sec. 5.6. have a long history, and a recent debate revolves around the claim that it can be derived from the other postulates of quantum theory (cf. Sec. 2.1). Many of these approaches have in common that their focus is on non-composite systems, either ignoring measurements on composite systems or making implicit assumptions about their consequences.
Our main tool will be the construction of foil theories of quantum theory (cf. Sec. 2.2), i.e. toy theories that help “to highlight the distinctive characteristics of quantum theory by contrasting with it” (adapted from [undefu]). Thus, we investigate the nature of the projection postulate by comparing quantum theory with similar theories that are equipped with different generalised state-update rules (GURs). In other words, we introduce generalised state-update theories (GUTs) for which the post-measurement states associated with a given outcome are not necessarily those prescribed by the Lüders rule. From the outset, we will take into account both single and composite systems. Alternative post-measurement rules have been considered occasionally (cf. Sec. 2.1), but we are not aware of any systematic approach.
@@ 2.1 The quantum-mechanical state update rule
Using quantum systems as information carriers has led to applications without classical equivalents. Quantum measurements, in particular, provide a resource that contributes, for example, to secure communication protocols [undefk], quantum teleportation [undefl] device-independent random number generators [undefv], and to speedy algorithms when using measurement-based quantum computation [undefo].
The advantages of quantum measurements arise from fundamental differences when compared with measurements in classical theories. Classical measurements are deterministic in that they reveal pre-existing values of observables, and the effect on the state of the observed system is negligible, at least in principle. In quantum theory, measurement outcomes are probabilistic, and the state of a system after a measurement coincides with the pre-measurement state only in exceptional cases. Phenomenologically, the quantum mechanical state update is governed by the projection postulate, which describes a “collapse” of the state to an eigenstate of the operator representing the observed outcome [undefac, undefaaz, undefaaf]. The collapse, also known as state reduction or Lüders rule [undefaaf], ensures deterministic repeatability: two measurements of an observable always yield identical outcomes when implemented in quick succession [undefw].
The standard postulates of quantum theory neither explain whether the state-update is a physical process [undefn, undefaa], nor do they define what constitutes a measurement [undefh, undefc]. The foundational questions resulting from this dilemma are often grouped under the term “measurement problem[undefr, undefaal]. The present work does not attempt to resolve the interpretational challenges posed by this long-standing problem.
Derivations of the Lüders rule have been given [undefi, undefas, undefaq, undefaad, undefaai] (see Appendix A for a survey), and the difference between von Neumann’s and Lüders’ formulations of the projection postulate has been discussed repeatedly [undefaaf, undefs, undefaay]. Some authors have argued that the projection postulate follows from the other standard postulates of quantum theory [undefaan, undefaak] and the debate of a recent derivation is ongoing [undefaaw, undefak, undefax, undefaaj, undefaav].
Occasionally, alternative post-measurement rules have been considered, either in the context of standard quantum theory [undefay, undefaz, undefaak] or of generalised probabilistic theories [undefaab]. A systematic investigation, as proposed here, has apparently not been carried out before.
@@ 2.2 Foil theories
Our approach is inspired by the constructive role of foil theories for the understanding of quantum theory [undefu]. The strategy in defining foil theories is to embed quantum theory in a larger set of theories by relaxing specific assumptions about the structure of the state space, say, or the form of the time evolution. We briefly describe typical examples of foil theories and point out that a systematic investigation of generalised state-update rules has, apparently, not yet been carried out.
@@
Generalised probabilistic theories (GPTs) [undefe, undefao, undefav], for example, represent a large and versatile collection of foil theories. They are defined by their own sets of states, observables, and rules for system composition, all of which may differ from those of standard quantum theory. GPTs have become a standard tool to explore and identify distinctive characteristics of quantum theory which they contain as a special case. Many GPTs share properties such as the no-cloning theorem, uncertainty relations, or teleportation with quantum theory [undefd, undefaam]. Some of them reproduce the standard quantum-mechanical correlations [undefaaac], while others exhibit superquantum correlations, e.g. Boxworld [undefaap].
@@
Ontological models describe a conceptually different class of foil theories not necessarily based on the familiar Hilbert space structure of quantum theory. They aim to explain its predictions in terms of an observer’s incomplete knowledge about the ontic state that provides an objective and complete description of the state of a system. In these models, non-classical features of quantum theory such as complementarity of observables, superdense coding, interference phenomena, or uncertainty relations may result from epistemic restrictions [undefaau, undeff].
@@
Other foil theories differ from quantum theory in a single key feature only, such as attempts to modify Born’s rule [undefal, undefaj], or the tensor product for system composition [undefad]. The consequences of associating observables with non-Hermitian operators have been investigated, using PT-symmetric [undefj] or normal [undefaas] operators, for example. Quantum-like theories defined over the real numbers [undefaax, undefaar] or quaternions [undefaf] rather than the field of complex numbers have also been studied.
@@
Nonlinear generalisations of Schrödinger’s equation have been suggested [undefm, undefa, undefae, undefay, undefaaq], most notably in Weinberg’s work [undefaaaa]. Gisin et al. [undefam, undefaat, undefg] pointed out that they imply violations of the no-signalling principle, although some specific nonlinear transformations were later found not to be ruled out by Gisin’s argument [undefx, undefay, undefae, undefar, undefaaq].
@@
In the specific context of modifications to the projection postulate, “causal quantum theory[undefaw] has been proposed, positing that the Lüders collapse is a well-defined physical process that satisfies strict local causality. Additionally, Kent introduced a new type of hypothetical measurement that is capable of revealing complete or partial information about the “local” state of a system, without actually inducing a state update [undefay, undefaz, undefax], an idea first presented in [undefp]. Such a measurement would, effectively, allow for yet other nonlinear state transformations without enabling superluminal signalling.
@@ 3.1 Axioms of quantum theory
    The probability that a measurement of $M$ performed on a system in the state $\rho$ yields the outcome represented by $P_{x}$ is given by the Born rule: $\text{prob}(x)=\text{Tr}\left[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}P_{x}}\,\rho\right]$.
@@
which sends $\rho$ to a sub-normalized state [^3] A sub-normalised state is a non-negative operator on a Hilbert space with trace less than or equal to $1$, i.e. $\rho\in\bar{\mathcal{S}}(\mathcal{H})=\{\lambda\rho:\lambda\in[0,1],\rho\in\mathcal{S}(\mathcal{H})\}$. in the space $\bar{\mathcal{S}}(\mathcal{H})$, and call it the Lüders rule, giving postulate ($\mathsf{L}$) its name. The trace of the post-measurement state, $\text{Tr}(\omega^{\mathsf{L}}(P_{x},\rho))=\text{Tr}(P_{x}\rho)$, yields the outcome probability as prescribed by the Born rule of postulate ($\mathsf{B}$). In addition, the maps (1) and (2) do not require the input state to be normalised. Replacing $\rho$ by $\lambda\rho$, $\lambda\in[0,1]$, shows that the Lüders rule is 1-homogeneous: $\omega^{\mathsf{L}}(P_{x},\lambda\rho)=\lambda\omega^{\mathsf{L}}(P_{x},\rho)$.
The alternative update rules $\omega(P_{x},\rho)$ considered in Sec. 5 will be modelled on the Lüders rule in the sense that they are 1-homogeneous, only depend on the measurement outcome $P_{x}$ and the initial state $\rho$, and agree with the Born rule for outcome probabilities. [^4] Postulate $(\mathsf{B})$ is understood to assign outcome probabilities to each individual measurement. Consequently, Born’s rule on its own does not account for correlations between outcomes of multiple measurements. Instead, the update rule is also responsible for the joint probabilities for the outcomes of measurements carried out sequentially, or by distinct parties in composite systems. This subtle point proves significant when examining modifications of the Lüders rule $(\mathsf{L})$; see Appendix A for further discussion. Further physically motivated properties that any acceptable update rule should satisfy will be discussed in Sec. 4.
@@ 3.2 Measurements on composite quantum systems
There are, however, consistency requirements when performing measurements on a composite system. If the outcome $P_{x}^{A}$ is non-degenerate, then the Lüders rule assigns a pure post-measurement state $\omega^{\mathsf{L}}(P_{x}^{A},\rho_{A})$ to subsystem $A$. Consequently, the post-measurement state of the composite system must be a product state, regardless of the initial state $\rho_{AB}$. On its own, the projection postulate ($\mathsf{L}$) provides, however, no information about the local state at $B$ upon performing a local measurement of $M_{A}$ on the constituent $A$. In principle, any density matrix $\rho_{B}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\text{Tr}_{A}}(\rho_{AB})\in\mathcal{S}(\mathcal{H}_{B})$ is compatible. For example, one could imagine that any local measurement at $A$ throws the state of the composite system into a product state, with the post-measurement state at $B$ being equal to the maximally mixed state $\rho_{B}\propto\mathbb{I}_{B}$.
@@ 4.2 Necessary properties of generalised update rules
The first three assumptions do not explicitly rely on the composite structure. Completeness requires that every conceivable measurement outcome specifies a unique post-measurement state. Context-independence asserts that the observed outcome and the pre-measurement state alone determine the post-measurement state. It introduces a form of non-contextuality [undefaac] that is present in quantum theory—namely, that all operationally equivalent experimental runs (i.e. involving the same initial state and the same measurement outcome) yield identical post-measurement states. Mathematically, completeness and context-independence determine the co-domain and the domain of the update rule $\omega_{AB}$ for a composite system, respectively.
@@
Completeness &\& context-independence A complete and context-independent update rule maps any combination of a (local) outcome, $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, and a sub-normalised state, $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, to another sub-normalised state,
@@
A generalised state-update rule $\omega$ is a set of functions $\{\omega_{AB}\}$—where each function $\omega_{AB}$ is associated with a pair of Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$—satisfying the following properties: (i) consistency with the Born rule (Eq. (6)); (ii) 1-homogeneity (Eq. (7)); and (iii) assumptions A1-A6 (Eqs. (8)-(13)), e.g. $\omega_{AB}:\mathcal{P}(\mathcal{H}_{A})\times\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\to\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ etc.
@@
How do the update rules defined here compare with quantum instruments [undefap], i.e. collections $\{\mathcal{I}_{x}\}_{x}$ of quantum operations [^10] Quantum operations are linear, completely positive and trace non-increasing maps defined on the space $\mathcal{L}(\mathcal{H})$ of bounded operators acting on $\mathcal{H}$. Quantum instruments are usually defined as mappings from an outcome space $\left(X,\Sigma\right)$ to the set of quantum operations. However, since we will only deal with discrete observables, an instrument is completely determined by the finite set of operations $\{\mathcal{I}_{x}\}_{x}$. that describe the effect on states of quantum measurements? The main difference is that the state-update rules satisfying our definition are not required to be linear or completely positive. Consequently, the set of alternative state-update rules is larger than the set of quantum instruments.
@@ 4.3 Updating (im-)proper mixtures
For a pure joint state $\rho_{AB}$, Eq. (11) applies irrespective of $\rho_{AB}$ being a product state or being entangled. Accordingly, $\omega_{A}\left(P_{x}^{A},\rho_{A}\right)$ represents the post-measurement state of system $\mathcal{H}_{A}$ when $\rho_{A}$ is either a pure state or an improper mixture, i.e. the reduced density operator of an entangled state. In contrast, proper mixtures result from classical uncertainty about the “true” preparation of a system. A proper mixture can be thought of as a Gemenge (German for ‘mixture’), i.e. as a collection $\mathcal{G}=\{(p_{1},\rho_{1}),(p_{2},\rho_{2}),\dots,(p_{n},\rho_{n})\}$ where $p_{i}\geq 0$, $\sum_{i}p_{i}=1$ and $\rho_{i}\in\bar{\mathcal{S}}(\mathcal{H})$ represent pure states or improper mixed states. The Gemenge $\mathcal{G}$ describes a system that is prepared in state $\rho_{i}$ with probability $p_{i}$ [undefr]. To any Gemenge there corresponds a unique density operator, $\rho_{\mathcal{G}}=\sum_{i=1}^{n}p_{i}\rho_{i}$.
@@
The equations (8)-(13) formalising assumptions A1-6 were formulated with pure and improper mixed states in mind. Nevertheless, they also ensure a consistent treatment of measurements on proper mixtures. For instance, the completeness A1 of $\omega$ guarantees that $\rho_{\mathcal{G}_{x}}$ in (14) is a valid state of the system. Similar arguments apply to assumptions A3-6, but not to context-independence A2, which requires minor adaptation. In fact, the post-measurement state $\rho_{\mathcal{G}_{x}}$ generally does not depend on the initial mixed state $\rho_{\mathcal{G}}$, but rather on the initial Gemenge $\mathcal{G}$. When proper mixtures are included, context-independence is generalised as follows: the observed outcome and the pre-measurement Gemenge alone determine the post-measurement state. [^12] If $\rho_{A}$ denotes either a pure or an improper mixed state, we can associate the trivial Gemenge $\mathcal{G}=\{(1,\rho_{A})\}$, describing a system prepared in the state $\rho_{A}$ with probability $1$.